# Language-Guided Self-Supervised Video Summarization Using Text Semantic Matching Considering the Diversity of the Video

### v1

@@ 4.7. Personalized Video Summarization
Figure 7. Visualization results of the personalized summarized video 6 in SumMe (“Car railcrossing”) generated by our model, where our model was guided to generate text summaries with a focus on the car after the accident by the LLMs.
We generate video summaries by calculating the similarity between the captions and the text summary produced by LLMs. This process allows us to influence the resulting video summary by adjusting the generated text summary according to user specifications. By controlling the content and focus of the text summary, we inherently affect the diversity and details emphasized in the video summary. For example, if you want to view footage of the car following an accident in the video 6 in SumMe, you can prompt the LLMs by stating, “I would like to watch the video that focus on the car after the accident.” Consequently, the model generates the video shown in Figure 7. In the generated video,a summarized version is generated that, as requested, focuses on the aftermath of the accident.
4.8. Limitation
The limitation of our approach is that our model does not consider the temporal relationships between input captions when generating video summaries. Although LLMs consider such temporal sequences when generating text summaries, our model calculates the similarity between the individual captions and the text summary independently. To address the temporal dependencies between input frames, some studies use a self-attention module for the entire video sequence as input (Zhong et al., 2022; Apostolidis et al., 2022; Abbasi and Saeedi, 2023). Additionally, BERT (Devlin et al., 2019) employs position embeddings to provide positional information within the input. Addressing the temporal dependencies of input captions when generating video summaries is a direction for our future work.
@@ 5. Conclusions
In this paper, we propose a novel LLM-guided self-supervised video summarization framework. To the best of our knowledge, this is the first work that leverage LLMs to address the video summarization task. We achieve frame-level scoring in the text semantic space. Additionally, we systematically analyze the characteristics of the datasets and construct a novel PDL function to create a more robust model tailored to the diversity of the video. The experimental results on the SumMe and TVSum datasets suggest that our proposed method produces superior quality summaries compared to many other SOTA unsupervised video summarization methods, demonstrating the effectiveness of our approach. Additionally, our method enables the creation of more personalized and customizable summaries tailored to the user’s objectives by allowing users to direct the generation of text summaries by LLMs. We hope that our proposed framework will inspire further advancements in the field of video summarization.
@@ Acknowledgements.
This work is partially and financially supported by the Beyond AI project of The University of Tokyo.
Appendix A Text summary generation
This supplementary material describes more details about the prompt design in generating text summary. We also show examples of the generated text summaries.
A.1. Prompt for text summary generation
Prompt for generating personalized text summary. In section 4.7 of the main draft, we mentioned that our proposed framework allows for the customization of the generated text summary according to the user query. Prompt 1 shows the detailed process of generating a personalized text summary. Specifically, we input the individually generated captions into [CAPTIONS] in chronological order, and the user query is fed into [USER QUERY].
Prompt 1: Our proposed prompt to generate text summary according to the user query.
You are an expert in video summarization.
In this task, you will create concise and personalized summaries from captions that have been created from video frames by an image captioning model.
Focus on extracting and highlighting only the elements of the captions that are directly relevant to the users specific interests.
Your goal is to emphasize the most significant details pertinent to the query while omitting any information that is not relevant, ensuring that the summary is insightful and precisely tailored to the users needs.
User query : ”””[USER QUERY]””””
Captions: ”””[CAPTIONS]”””
The final summary should be delivered in JSON format in a single line (~80 words), perfectly encapsulating the users query with accuracy and relevance, and avoiding any extraneous content.
A.2. Examples for generated text summary
@@
Generated general text summary. A general text summary example, generated by GPT-4 using Prompt LABEL:summary, is presented in Text Summary 1. It describes the scene where many vehicles are visible while the railroad crossing is closed, and the crowd that gathers following the accident. This text summary captures both the moments before and after the accident, making it a suitable summary for the video.

### v2

@@ 4.7. Personalized Video Summarization
Figure 8. Visualization result of the personalized summarized video 6 in SumMe (“Car railcrossing”) generated by our model, where our model was guided by the LLMs to generate text summaries with a focus on the car after the accident.
We generate video summaries by calculating the similarity between the captions and the text summary produced by LLMs. This process allows us to influence the resulting video summary by adjusting the generated text summary according to user specifications. By incorporating user queries as prompts into the LLMs, we can flexibly control the content and focus of the text summary, inherently affecting the diversity and details emphasized in the video summary. The idea of leveraging LLMs to allow users to specify what they want to see in the video bridges the semantic gap, making the framework flexible enough to summarize videos across various domains and effectively achieve personalization.
For example, if you want to view footage of the car following an accident in the video 6 in SumMe, you can prompt the LLMs by stating, “I would like to watch the video that focuses on the car after the accident.” Consequently, the model generates the video shown in Figure 8. In the generated video, a summarized version is generated that, as requested, focuses on the aftermath of the accident. The prompt design and more examples using the Mr.HiSum dataset (Sul et al., 2023) are provided in the supplementary material.
4.8. Limitations
The limitation of our approach is that our model does not account for the temporal dependencies between input captions although LLMs consider temporal sequences when generating text summaries. Additionally, the captions generated by the image captioning model do not always perfectly describe the frames. We will address these issues in our future work.
@@ 5. Conclusions
In this paper, we propose a novel LLM-guided self-supervised video summarization framework. Our method eliminates the need for extensive data annotation and reduces subjectivity. We achieve frame-level scoring in the text semantic space. Additionally, we systematically analyze the characteristics of the datasets and mathematically define the diversity of the video. Subsequently, we construct a novel PDL function to create a more robust model tailored to the diversity of the video. The experimental results suggest that our proposed method achieves SOTA performance on the SumMe dataset and the second-best results on the TVSum dataset, demonstrating the effectiveness of our approach. Additionally, our proposed framework flexibly enables the creation of personalized and customizable summaries tailored to the user’s objectives by allowing users to direct the generation of text summaries by LLMs. This paper paves a new way for video summarization and is crucial for real-world scenarios where the video text description is not always available. We hope our framework will inspire further advancements in the field of video summarization.
@@ Acknowledgements.
This work is partially financially supported by the Beyond AI project of The University of Tokyo.
This supplementary material provides the implementation details and further details of the personalized video summarization.
Appendix A Implementation Details
A.1. Comparison of caption diversity
We generate individual captions from downsampled frames using an image captioning model. One main reason for using an image captioning model instead of a video captioning model for caption generation is that using an image captioning model allows for more accurate assessment of the content of each frame and calculation of frame-level scores. Regarding the prompt, we compare three prompts: “a scene of”, “a frame of”, and “a photo of”, using two image captioning models: the Generative Image-to-text transformer (GIT) (Wang et al., 2022) and the Bootstrapping Language-Image Pre-training with frozen unimodal models 2 (BLIP-2) (Li et al., 2023b). The average percentages of generated unique captions within each video are shown in Table 3. Comparing the three prompts, “a photo of” results in the highest uniqueness in captions across both datasets and both image captioning models. When comparing the two image captioning models, GIT (Wang et al., 2022) demonstrates higher diversity compared to BLIP-2 (Li et al., 2023b), indicating minimal duplication and a lower occurrence of repetitive expressions. Therefore, we finally use GIT as the image captioning model and “a photo of” as the prompt to generate descriptive captions from individual downsampled video frames, as this combination results in more diverse captions. This diversity allows for a more accurate capture of the dynamic actions within the video, providing a richer and more detailed representation of the content.
A.2. Experiment Details
In the experiment investigating the contributions of the regularization loss in Section 3.3 of the main draft, we train our model with a learning rate of $1\times 10^{-5}$, and set the margin $m$ in Eq. (3) to 0.15 for both the SumMe and TVSum datasets. For the general video summarization in Section 4.2 of the main draft, we set the learning rate of $5\times 10^{-5}$ and $1\times 10^{-5}$, the margin $m$ in Eq. (3) to 0.11 and 0.06 for the SumMe and TVSum datasets, respectively.
A.3. Prompt for text summary generation
Prompt for generating personalized text summary. In section 4.7 of the main draft, we mentioned that our proposed framework allows for the customization of the generated text summary according to the user query. Prompt 1 shows the detailed process of generating a personalized text summary. Specifically, we input the individually generated captions into [CAPTIONS] in chronological order, and the user query is fed into [USER QUERY]. Prompt 1: Our proposed prompt to generate text summary according to the user query. You are an expert in video summarization. In this task, you will create concise and personalized summaries from captions that have been created from video frames by an image captioning model. Focus on extracting and highlighting only the elements of the captions that are directly relevant to the users specific interests. Your goal is to emphasize the most significant details pertinent to the query while omitting any information that is not relevant, ensuring that the summary is insightful and precisely tailored to the users needs. User query : ”””[USER QUERY]”””” Captions: ”””[CAPTIONS]””” The final summary should be delivered in JSON format in a single line (~80 words), perfectly encapsulating the users query with accuracy and relevance, and avoiding any extraneous content.
Table 3. Comparison of the average percentages of generated unique captions on the SumMe and TVSum datasets. The bolded items represent the best results.
Prompt SumMe (Gygli et al., 2014) TVSum (Song et al., 2015)
BLIP-2 (Li et al., 2023b) GIT (Wang et al., 2022) BLIP-2 (Li et al., 2023b) GIT (Wang et al., 2022)
“a scene of” 28 69 36 76
“a frame of” 28 68 36 76
“a photo of” 29 69 37 76
A.4. Examples for generated text summary
@@
Generated general text summary. A general text summary example, generated by GPT-4 using the Chain of Density prompt, is presented in Text Summary 1. It describes the scene in which many vehicles are visible while the railroad crossing is closed, and the crowd that gathers following the accident. This text summary captures both the moments before and after the accident, making it a suitable summary for the video.
@@
Appendix B Personalized video summarization
B.1. Dataset
In this section, we present more details of the personalized video summarization using the videos in the Mr.HiSum (Sul et al., 2023) dataset. Mr.HiSum consists of YouTube videos within the YouTube-8M dataset (Abu-El-Haija et al., 2016), specifically those that have highlight labels available on YouTube.
B.2. Implementation Details
The videos in the Mr.HiSum dataset are downsampled into one frame per second (fps) and scene boundaries are obtained using the KTS algorithm (Potapov et al., 2014). Therefore, we downsample the video into one fps. The other experimental settings are the same as Section 4.7 of the main draft. We specifically use the movie trailers titled “Restless — Official Trailer HD (2011)”  [^2] https://www.youtube.com/watch?v=pzUKbPAynxU and “At Any Price Official Trailer”  [^3] https://www.youtube.com/watch?v=P_L35FyXqTA in the Mr.HiSum dataset. The original videos, named “Restless_original.mp4” and “At_Any_Price_original.mp4”, are available in the folder linked on Google Drive  [^4] https://drive.google.com/drive/folders/1HPibyBmmEqWQGAeMJaRR2nJ0vbAY7nBs?usp=sharing. For each video, we use two different user queries, providing two personalized video summaries per video. The generated personalized videos are also provided as MP4 files 4.
B.3. Evaluation Metrics
In order to quantitatively evaluate the generated personalized video summaries, we use recall and precision. Specifically, given the frames in the generated personalized video summary $\rm{P}$ and the ground truth frames that correspond to the user query $\rm{Q}$, the precision $\rm{p}$ and the recall $\rm{r}$ are defined as follows:
(12) $\rm{p}=\frac{{\rm{P}}\cap{\rm{Q}}}{\text{Len}({\rm{P}})},\quad{\rm{r}}=\frac{{\rm{P}}\cap{\rm{Q}}}{\text{Len}(\rm{Q})}.\quad$
B.4. Generated text summaries
The original video titled “Restless — Official Trailer HD (2011)” depicts a romantic story between a man and a woman. The man meets the woman by chance at a funeral, and as they get to know each other, their love gradually grows. Text Summary 3 and 4 show the results for the user queries “I would like to watch a video that focuses on the background scenery and landscapes.” and “I would like to watch a video that focuses on the conversations and interactions between characters.”, respectively.
Text Summary 3: Generated personalized text summary for “Restless — Official Trailer HD (2011)” with a focus on the background scenery and landscapes.
The video features diverse backgrounds, including a man and a woman standing by a river with a red flag, highlighting a picturesque river scene. Two people sit on a bench next to a lake under an orange sky, offering a serene lakeside view. A couple sits on a bench by a lake, looking at trees and the water, emphasizing peaceful lake scenery. These scenes collectively showcase the videos focus on beautiful landscapes and tranquil outdoor settings.
Text Summary 4: Generated personalized text summary for “Restless — Official Trailer HD (2011)” with a focus on the conversations and interactions between characters.
The video features various interactions among characters, including a man and woman sitting at a dinner table, a couple discussing by a river, individuals conversing in a park, a woman talking to a man in a kitchen, and a couple kissing in the dark. It highlights personal and intimate moments, such as a boy talking to a girl, a couple walking and holding hands, and two people kissing on a sidewalk, emphasizing the focus on conversations and interactions between characters.
The original video titled “At Any Price Official Trailer” depicts the story of a famous farmer’s son who defies his father’s wishes and pursues his dream of becoming a professional race car driver instead of taking over the family business. In addition to the movie trailer, this video includes footage of a female speaker speaking at the beginning and end, which together make up roughly a quarter of the total runtime. Based on this content, we provide two types of user queries and evaluate whether the generated personalized video summaries correspond to the user queries. Text Summary 5 and 6 show the results for the user queries “I would like to watch a video that focuses on scenes related to cars.” and “I would like to watch a video that focuses on rural life and scenery.”, respectively. Text Summary 5: Generated personalized text summary for “At Any Price Official Trailer” with a focus on the car related scenes. The video features various scenes related to cars, including a man and woman watching an orange race car, a woman waving to the crowd with a car on track, posters mentioning race car’, scenes of a race car with the number 25, a group walking in front of a race car numbered 48, a car with clio on the bottom in the middle of a race track, a couple kissing in a car with clio visible, and a car driving down a dirt road kicking up dust.”
Text Summary 6: Generated personalized text summary for “At Any Price Official Trailer” with a focus on the rural life and scenery.
The video captures the essence of rural life and scenery, featuring a man in a field with a tractor and a large building in the background, a green tractor parked in front of a silo, and a farm with a silo and a red tractor. It also shows a man and woman looking out a window at a corn field, a red grain silo in front of a barn, and a car driving down a dirt road next to a field of corn, highlighting the agricultural and serene aspects of rural living.”
B.5. Visualizations and Results
((a))
((b))
((c))
Figure 9. Visualization of the video titled “Restless — Official Trailer HD (2011)” focusing on the background scenery and landscapes.
((a))
((b))
((c))
Figure 10. Visualization of the video titled “Restless — Official Trailer HD (2011)” focusing on the conversations and interactions between characters.
Figure 9(a), 10(a), 11(a), and  12(a) show the ground truth frames we choose that are related to the user query. It should be noted that the video is downsampled to one fps. Additionally, if a frame corresponding to the user query exists within each scene, one representative frame is selected. Also, Figure 9(b), 10(b), 11(b), and  12(b) show the downsampled frames of our generated personalized video, which has also been downsampled to one fps. Additionally, Figure 9(c), 10(c), 11(c), and  12(c) show the correct frames generated by our model.
Figure 9 and 10 show the visualizations of the video titled “Restless — Official Trailer HD (2011)” when the user queries are “I would like to watch a video that focuses on the background scenery and landscapes.” and “I would like to watch a video that focuses on the conversations and interactions between characters.”, respectively. The generated personalized videos, named “Restless_scenery.mp4” and “Restless_conversation.mp4”, are available in the folder linked on Google Drive 4. Figure 9(b) shows that out of the 14 frames in the generated personalized video, 10 frames correspond to the user query focusing on the background scenery and landscapes. Therefore, the precision is 0.714. Also, Figure 9(c) shows that, out of the 8 ground truth frames, our model selects 5 frames. Therefore, the recall is 0.625. In the same way, Figure 10 shows that out of the 14 frames in the generated personalized video, 13 frames correspond to the user query focusing on the conversations and interactions between characters (${\rm{p}}=0.929$). Additionally, out of 17 ground truth frames, our model selects 8 frames (${\rm{r}}=0.471$). These results indicate that, while some ground truth scenes are not selected due to the limitations of the summarized video length, the majority of the scenes within the generated personalized summaries correspond to the user query.
((a))
((b))
((c))
Figure 11. Visualization of the video titled “At Any Price Official Trailer” focusing on the car related scenes.
((a))
((b))
((c))
Figure 12. Visualization of the video titled “At Any Price Official Trailer” focusing on the rural life and scenery.
Figure 11 and 12 show the visualizations of the video titled “At Any Price Official Trailer” for the user queries “I would like to watch a video that focuses on scenes related to cars.” and “I would like to watch a video that focuses on rural life and scenery.”, respectively. The generated personalized videos, named “At_Any_Price_car.mp4” and “At_Any_Price_rural.mp4”, are available in the Google Drive link 4. Figure 11(b) shows that out of the 22 frames in the generated personalized video, 15 frames correspond to the user query focusing on the car related scenes (${\rm{p}}=0.682$). Also, Figure 11(c) shows that, out of the 16 ground truth frames, our model selects 11 frames (${\rm{r}}=0.688$). Similarly, Figure 10 shows that out of the 23 frames in the generated personalized video, 15 frames correspond to the user query focusing on the rural life and scenery (${\rm{p}}=0.652$). Additionally, out of 15 ground truth frames, our model selects 9 frames (${\rm{r}}=0.600$). These results demonstrate that the generated personalized video summaries successfully reflect distinctly different user queries, such as those related to cars and those depicting rural life and scenery, by creating personalized text summaries tailored to each user query.
B.6. Limitations
The limitation of our personalized video summarization approach is that we did not perform parameter searching for the Mr.HiSum dataset, so we believe further optimization of the parameter settings can ensure better results. Additionally, we solve the knapsack problem to extract a subset of scenes. Among the scenes corresponding to the user query, some scenes corresponding to the user query are not selected due to their long length, despite having high scene-level scores. This issue arises from solving the knapsack problem to create the summary, as mentioned in Section 4.3 of the main draft.